OpenStudy (meilendurcer):

-1/4 + (-2/7) A. -15/28 B. -3/11 C. 9/14 D. 3/28

OpenStudy (meilendurcer):

@hartnn can you help?

hartnn (hartnn):

-1/4-2/7 we need to make the denominator same. what is LCM of 4 and 7 ?

OpenStudy (meilendurcer):

4 and 7? LCM

hartnn (hartnn):

yes, LCM. what is the Least Common Multiple of 4 and 7

OpenStudy (meilendurcer):

oh i thought their lcm's were 4 and 7

hartnn (hartnn):

nopes, common multiple is 28 ..... so multiply and divide 1st fraction by 7 and multiply and divide 2nd fraction by 4

OpenStudy (meilendurcer):

oh.... ok yea, 7/-8?

hartnn (hartnn):

O.o how? \(\huge \frac{-1}{4}\times \frac{7}{7} -\frac{2}{7}\times \frac{4}{4}=\frac{-7-8}{28}=?\)

OpenStudy (meilendurcer):

-7.28? im sorry i am doing this on calculator

hartnn (hartnn):

-7-8 = -15 no need of calculator -15/28 is in your options

OpenStudy (meilendurcer):

omg thank you so much!

OpenStudy (meilendurcer):

@hartnn this one also? 11 1/3 ÷ 4 1/4?

hartnn (hartnn):

11 1/3 = ?